OpenStudy (anonymous):

Sandra plans to fill a corner of her yard with decorative stones. Each row has four more stones than the prior row. If there were five stones in the first row, how many stones were used to fill all 44 rows?

OpenStudy (anonymous):

\[S_{44} = \frac n2(2a_1 + (n - 1)d\]where: n = 44 \(a_1\) = 5 d = 4 Can you use the formula to figure it out?

OpenStudy (anonymous):

I'm unsure i can try though 44/2(2(5)+(44-1)4 22(10+43)4 22(53)4 1166(4) =4664

OpenStudy (anonymous):

@Calcmathlete is that correct ?

OpenStudy (anonymous):

It should be 4004. 22(10 + (43)4) = 4004 See?

OpenStudy (anonymous):

okay thanks